# Is the Faber-Krahn inequality true for the Stokes operator?

### v1

@@ 5 The ball does not minimise $F$ in dimension 3: proof of Theorem 3
$\lambda_{1}({\mathbb{B}}_{3})$ has multiplicity 3. An orthogonal basis of the associated eigenspace $E_{1}$ is given by
@@
$\int_{{\mathbb{B}}_{3}}\|{\mathbf{U}}_{i}\|^{2}=\frac{2\pi}{3\omega^{4}}\left(\omega^{2}-\sin^{2}\omega\right)=:A.$
@@
Let us now give another proof of Theorem 3 that only relies on the analysis of the semi-differential as given section 4. Recall that the semi-differential of $\lambda_{1}$ at $B_{3}$ in direction $\Phi\in\mathscr{X}$ (see Lemma 3 for the definition of $\mathscr{X}$) reads
@@
$\nabla\mathbf{U}_{1}\nu=\begin{pmatrix}0\\ -z\psi^{\prime}(1)\\ y\psi^{\prime}(1)\end{pmatrix},\quad\nabla\mathbf{U}_{2}\nu=\begin{pmatrix}z\psi^{\prime}(1)\\ 0\\ -x\psi^{\prime}(1)\end{pmatrix},\quad\nabla\mathbf{U}_{3}\nu=\begin{pmatrix}-y\psi^{\prime}(1)\\ x\psi^{\prime}(1)\\ 0\end{pmatrix}$
@@
$\mathcal{M}_{\partial{\mathbb{B}}_{3}}(\langle\Phi,\nu\rangle)=\psi^{\prime}(1)^{2}\begin{pmatrix}-\int_{\partial B_{3}}(y^{2}+z^{2})\langle\Phi,\nu\rangle&\int_{\partial B_{3}}xy\langle\Phi,\nu\rangle&\int_{\partial B_{3}}xz\langle\Phi,\nu\rangle\\ \int_{\partial B_{3}}xy\langle\Phi,\nu\rangle&-\int_{\partial B_{3}}(x^{2}+z^{2})\langle\Phi,\nu\rangle&\int_{\partial B_{3}}yz\langle\Phi,\nu\rangle\\ \int_{\partial B_{3}}xz\langle\Phi,\nu\rangle&\int_{\partial B_{3}}yz\langle\Phi,\nu\rangle&-\int_{\partial B_{3}}(x^{2}+y^{2})\langle\Phi,\nu\rangle\\ \end{pmatrix}.$
@@
$=$ $|{\mathbb{B}}_{3}|^{-1/3}\left(\frac{2}{3}\omega^{2}\int_{\partial{\mathbb{B}}_{3}}\langle\Phi,\nu\rangle+\omega^{2}\mu_{1}\left(\mathcal{M}_{\partial{\mathbb{B}}_{3}}(\langle\Phi,\nu\rangle)\right)\right)$
$=$ $|{\mathbb{B}}_{3}|^{-1/3}\mu_{1}\left(\frac{2}{3}\omega^{2}\int_{\partial{\mathbb{B}}_{3}}\langle\Phi,\nu\rangle\operatorname{I}_{3}+\omega^{2}\mathcal{M}_{\partial{\mathbb{B}}_{3}}(\langle\Phi,\nu\rangle)\right)$
where $\operatorname{I}_{3}$ denotes the identity matrix of size 3. Introduce the $3\times 3$ matrix
$\widehat{M}_{\Phi}:=\frac{2}{3}\omega^{2}\int_{\partial{\mathbb{B}}_{3}}\langle\Phi,\nu\rangle\operatorname{I}_{3}+\omega^{2}\mathcal{M}_{\partial{\mathbb{B}}_{3}}(\langle\Phi,\nu\rangle)$
@@
$\operatorname{Tr}\widehat{M}_{\Phi}=2\omega^{2}-2\frac{\omega^{4}\sin^{2}\omega}{\omega^{2}-\sin^{2}\omega}=0,$
@@
$\int_{\partial B_{3}}yz\langle\Phi,\nu\rangle\neq 0.$
@@ 6 Conclusion and open problems
Although we have proved that, in $\mathbb{R}^{2}$, the ball ${\mathbb{B}}_{2}$ is likely to be an optimiser, it is not clear to us how one would approach such a result. Indeed, in the vectorial case, standard rearrangement and symmetrisation tools are bound to fail. One might then try to use an alternative approach to the Faber-Krahn inequality, typically using an overdetermined problem framework, but, to the best of our knowledge, no generalisation of the Serrin theorem can accommodate incompressibility constraints. In three dimensions, it seems useful to perform some numerical simulations to have at least an idea of the possible optimal sets. Following our Corollary 1, one can wonder whether it is a particular torus;
@@
Following Theorem 1, we know that an optimal domain ${\Omega}^{*}$ exists in the class of quasi-open sets. Nevertheless, it seems extremely ambitious, at this stage, to develop an a priori regularity theory for this optimal set. Following the usual approach to such problems, see e.g. [28, chapter3], the first step in that direction would be to establish that ${\Omega}^{*}$ is, in fact, open. To derive further properties, the optimality conditions given in Theorem 4 would probably play a crucial role in the development of an appropriate blow-up theory.
@@ Acknowledgments
The authors warmly thank Dorin Bucur, who provided the main argument in the proof of existence. I. M-F thanks Davide Buoso for helpful conversations. I. M-F was partially funded by a PSL Young Researcher Starting Grant 2023. The support of the CNRS through a 2023 PEPS project is also acknowledged.
@@ Appendix A The first Stokes eigenvalue in ${\mathbb{B}}_{3}$
$\lambda_{1}({\mathbb{B}}_{3})$ has multiplicity 3. An orthonormal basis of the associated eigenspace $E_{1}$ is given by
@@
To study the existence of such functions, we can follow the approach of [44, Lemma 2]. Dropping the indices $i$ in ${\mathbf{u}}_{i,\pm}$ and $v_{i}$, the function ${\mathbf{u}}_{\pm}=\pm\frac{v}{r}{\vec{e}_{r}}+w{\vec{e}_{\theta}}+z{\vec{e}_{\phi}}$ must satisfy
@@
$=$ $\frac{8\pi}{3\omega^{3}}\left(\frac{2\cos^{2}\omega+\omega^{2}+\omega\cos\omega\sin\omega-2}{2\omega^{2}}\right).$
The desired expression follows by using that $\tan\omega=\omega$.

### v2

@@ 5 The ball does not minimise $F$ in dimension 3: proof of Theorem 3
$\lambda_{1}({\mathbb{B}}_{3})$ is equal to $\omega^{2}$ and has multiplicity 3. An orthogonal basis of the associated eigenspace $E_{1}$ is given by
@@
$\int_{{\mathbb{B}}_{3}}\|{\mathbf{U}}_{i}\|^{2}=\frac{4\pi}{3\omega^{4}}\left(\omega^{2}-\sin^{2}\omega\right)=\frac{4\pi}{3}\cos^{2}\omega=:A^{2}.$
@@
Let us now give another proof of Theorem 3 that only relies on the analysis of the semi-differential as given in section 4. Recall that the semi-differential of $\lambda_{1}$ at ${\mathbb{B}}_{3}\color[rgb]{0,0,0}$ in direction $\Phi\in\mathscr{X}$ (see Lemma 3 for the definition of $\mathscr{X}$) reads
@@
$\nabla\mathbf{U}_{1}\nu=\begin{pmatrix}0\\ z\psi^{\prime}(1)\\ -y\psi^{\prime}(1)\end{pmatrix},\quad\nabla\mathbf{U}_{2}\nu=\begin{pmatrix}-z\psi^{\prime}(1)\\ 0\\ x\psi^{\prime}(1)\end{pmatrix},\quad\nabla\mathbf{U}_{3}\nu=\begin{pmatrix}-y\psi^{\prime}(1)\\ x\psi^{\prime}(1)\\ 0\end{pmatrix}$
@@
${\mathcal{M}_{\partial{\mathbb{B}}_{3}}(\langle\Phi,\nu\rangle)=\psi^{\prime}(1)^{2}\begin{pmatrix}-\int_{\partial{\mathbb{B}}_{3}}(y^{2}+z^{2})\langle\Phi,\nu\rangle&\int_{\partial{\mathbb{B}}_{3}}xy\langle\Phi,\nu\rangle&-\int_{\partial{\mathbb{B}}_{3}}xz\langle\Phi,\nu\rangle\\ \int_{\partial{\mathbb{B}}_{3}}xy\langle\Phi,\nu\rangle&-\int_{\partial{\mathbb{B}}_{3}}(x^{2}+z^{2})\langle\Phi,\nu\rangle&-\int_{\partial{\mathbb{B}}_{3}}yz\langle\Phi,\nu\rangle\\ -\int_{\partial{\mathbb{B}}_{3}}xz\langle\Phi,\nu\rangle&-\int_{\partial{\mathbb{B}}_{3}}yz\langle\Phi,\nu\rangle&-\int_{\partial{\mathbb{B}}_{3}}(x^{2}+y^{2})\langle\Phi,\nu\rangle\\ \end{pmatrix}.}$
@@
Moreover, since
$\frac{|{\mathbb{B}}_{3}|}{A^{2}}=\frac{4\pi}{3}\cdot\frac{3\omega^{2}}{4\pi\sin^{2}\omega}=\frac{\omega^{2}}{\sin^{2}\omega},$
it follows that
$\langle\partial\mathcal{F}({\mathbb{B}}_{3}),\Phi\rangle$ $=$ $\omega^{2}|{\mathbb{B}}_{3}|^{-1/3}\left(\frac{2}{3}\int_{\partial{\mathbb{B}}_{3}}\langle\Phi,\nu\rangle+\frac{1}{\sin^{2}\omega}\mu_{1}\left(\mathcal{M}_{\partial{\mathbb{B}}_{3}}(\langle\Phi,\nu\rangle)\right)\right)$
$=$ $\omega^{2}|{\mathbb{B}}_{3}|^{-1/3}\mu_{1}\left(\widehat{M}_{\Phi}\right)$
where $\widehat{M}_{\Phi}$ denotes the $3\times 3$ matrix
$\widehat{M}_{\Phi}:=\frac{2}{3}\int_{\partial{\mathbb{B}}_{3}}\langle\Phi,\nu\rangle\operatorname{I}_{3}+\frac{1}{\sin^{2}\omega}\mathcal{M}_{\partial{\mathbb{B}}_{3}}(\langle\Phi,\nu\rangle),$
$\operatorname{I}_{3}$ is the identity matrix of size 3 and $\mu_{1}\left(\widehat{M}_{\Phi}\right)$ denotes the lowest eigenvalue of the matrix $\widehat{M}_{\Phi}$.
@@
$\operatorname{Tr}\widehat{M}_{\Phi}=\left(2-2\frac{\sin^{2}\omega}{\sin^{2}\omega}\right)\int_{\partial{\mathbb{B}}_{3}}\langle\Phi,\nu\rangle=0,$
@@
$\int_{\partial{\mathbb{B}}_{3}}yz\langle\Phi,\nu\rangle\neq 0.$
@@ 6 Conclusion and open problems
Although we have proved that, in $\mathbb{R}^{2}$, the ball ${\mathbb{B}}_{2}$ is likely to be an optimiser, it is not clear to us how one would approach such a result. As detailed in Remark 3, if we can prove that there exists a simply connected and regular minimiser, then it is necessarily a disc. The results in our paper complete the intuition that the disc minimises the Stokes eigenvalue under volume constraints in dimension 2. However, the proof of this result is an open problem as far as we know.
Indeed, in the vectorial case, standard rearrangement and symmetrisation tools are bound to fail. One might then try to use an alternative approach to the Faber-Krahn inequality, typically using an overdetermined problem framework, but, to the best of our knowledge, no generalisation of the Serrin theorem can accommodate incompressibility constraints. In three dimensions, it seems useful to perform some numerical simulations to have at least an idea of the possible optimal sets. Following our Corollary 1, one can wonder whether it is a particular torus;
@@
Following Theorem 1, we know that an optimal domain ${\Omega}^{*}$ exists in the class of quasi-open sets. Nevertheless, it seems extremely ambitious, at this stage, to develop an a priori regularity theory for this optimal set. Following the usual approach to such problems, see e.g. [32, chapter3], the first step in that direction would be to establish that ${\Omega}^{*}$ is, in fact, open. To derive further properties, the optimality conditions given in Theorem 4 would probably play a crucial role in the development of an appropriate blow-up theory.
@@ Acknowledgments
The authors warmly thank Dorin Bucur, who provided the main argument in the proof of existence. The authors also thank Davide Buoso, Wadim Gerner and Antoine Lemenant for helpful conversations and their comments on the results of this article. Finally, the authors wish to warmly thank the anonymous reviewer, whose comments helped improve and clarify this article.
I. M-F was partially funded by a PSL Young Researcher Starting Grant 2023. The support of the CNRS through a 2023 PEPS project is also acknowledged.
Data Availability Statement
No datasets were generated or analysed during the current study.
@@ Appendix A The first Stokes eigenvalue in ${\mathbb{B}}_{3}$
Let $\omega>0$ be the first positive root of $tan(x)=x$. Then $\lambda_{1}({\mathbb{B}}_{3})$ is equal to $\omega^{2}$ and has multiplicity 3. An orthogonal basis of the associated eigenspace $E_{1}$ is given by
@@
Finally, for any $i\in\{1,2,3\}$, there holds
$\int_{{\mathbb{B}}_{3}}\|{\mathbf{U}}_{i}\|^{2}=\frac{4\pi}{3\omega^{4}}\left(\omega^{2}-\sin^{2}\omega\right)=\frac{4\pi}{3}\cdot\frac{\sin^{2}\omega}{\omega^{2}}=:A^{2}.$
@@
To study the existence of such functions, we can follow the approach of [51, Lemma 2]. Dropping the indices $i$ in ${\mathbf{u}}_{i,\pm}$ and $v_{i}$, the function ${\mathbf{u}}_{\pm}=\pm\frac{v}{r}{\vec{e}_{r}}+w{\vec{e}_{\theta}}+z{\vec{e}_{\phi}}$ must satisfy
@@
$=$ $\frac{8\pi}{3}\left(\frac{2\cos^{2}\omega+\omega^{2}+\omega\cos\omega\sin\omega-2}{2\omega^{4}}\right).$
The desired expression follows easily by using that $\omega\cos\omega=\sin\omega$.
Appendix B Existence of an optimal shape under a box constraint
In this appendix, we prove the following result:
Proposition 6.
Let $D\subset\mathbb{R}^{d}$ be a fixed compact domain of $\mathbb{R}^{d}$ and define the admissible class
$\widehat{\mathscr{O}}_{V_{0}}=$Ω quasi-open included in $\left\{\Omega\textrm{ quasi-open included in }D\,,|\Omega|\leqslant V_{0}\right\}.$
The shape optimisation problem
$\inf_{{\Omega}\in\widehat{\mathscr{O}}_{V_{0}}}\lambda({\Omega})$ (41)
has a solution.
Proof of Proposition 6.
We introduce, for $\Omega\in\widehat{\mathcal{O}}_{V_{0}}$, the quantity $\widehat{\lambda_{1}}(V_{0}\color[rgb]{0,0,0})$ given by
$\widehat{\lambda_{1}}(V_{0})=\inf_{\begin{subarray}{c}\mathbf{u}\in W^{1,2}_{0}(D;\mathbb{R}^{3})\\ \int_{D}\mathbf{u}\cdot\nabla v=0\ \forall v\in H^{1}(D)\\ |\Omega_{\mathbf{u}}|\leqslant V_{0}\\ \int_{\Omega_{\mathbf{u}}}\|{\mathbf{u}}\|^{2}=1\end{subarray}}{\int_{D}\|\nabla\mathbf{u}\|^{2}}$ (42)
where $\Omega_{\mathbf{u}}$ denotes the quasi open set $\{\mathbf{u}\neq 0\}$.
Let $\{{\mathbf{u}}_{k}\}_{k\in\mathbb{N}}$ be minimising for Problem (42). For any $k\in\mathbb{N}$, let $\Omega_{k}:=\{{\mathbf{u}}_{k}\neq 0\}$. In view of the Poincaré inequality [34, Lemma 4.5.3], $\{{\mathbf{u}}_{k}\}_{k\in\mathbb{N}}$ is bounded in $W^{1,2}_{0}(D;\mathbb{R}^{3})$. It follows that there exists $\mathbf{u}^{\star}\in W^{1,2}_{0}(D;\mathbb{R}^{3})$ such that $\{{\mathbf{u}}_{k}\}_{k\in\mathbb{N}}$ converges to $\mathbf{u}^{\star}$ up to a (not relabelled) subsequence, weakly in $W^{1,2}_{0}$ and strongly in $L^{2}$. Thus, $\{u_{k}\}_{k\in\mathbb{N}}$ converges almost everywhere to $\mathbf{u}^{\star}$, and therefore
$|\Omega_{\mathbf{u^{\star}}}|\leqslant\liminf_{k\to+\infty}|\Omega_{k}|\leqslant V_{0}$
where $\Omega_{\mathbf{u^{\star}}}$ denotes the quasi-open set $\{\mathbf{u^{\star}}\neq 0\}$. We infer that $\Omega_{\mathbf{u^{\star}}}$ belongs to $\widehat{\mathscr{O}}_{V_{0}}$. Furthermore, since $\mathbf{u^{\star}}=0$ quasi-everywhere on $D\backslash\Omega_{\mathbf{u}^{\star}}$, by weak $H^{1}$-convergence, one has
by minimality, whence the equality of these quantities. By strong convergence in $L^{2}(D;\mathbb{R}^{3})$ and weak convergence in $W^{1,2}_{0}(D;\mathbb{R}^{3})$, there holds
$\int_{D}\|\mathbf{u}^{*}\|^{2}=1$
and, for any $v\in W^{1,2}(D)$,
Thus, ${\Omega}_{{\mathbf{u}}^{*}}$ is a solution of (42). The proof is now concluded. ∎